# Change-Agent: Towards Interactive Comprehensive Remote Sensing Change Interpretation and Analysis

### v2

@@ Abstract
Monitoring changes in the Earth’s surface is crucial for understanding natural processes and human impacts, necessitating precise and comprehensive interpretation methodologies. Remote sensing satellite imagery offers a unique perspective for monitoring these changes, leading to the emergence of remote sensing image change interpretation (RSICI) as a significant research focus. Current RSICI technology encompasses change detection and change captioning, each with its limitations in providing comprehensive interpretation. To address this, we propose an interactive Change-Agent, which can follow user instructions to achieve comprehensive change interpretation and insightful analysis according to user instructions, such as change detection and change captioning, change object counting, change cause analysis, etc. The Change-Agent integrates a multi-level change interpretation (MCI) model as the eyes and a large language model (LLM) as the brain. The MCI model contains two branches of pixel-level change detection and semantic-level change captioning, in which multiple BI-temporal Iterative Interaction (BI3) layers utilize Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) modules to enhance the model’s discriminative feature representation capabilities. To support the training of the MCI model, we build the LEVIR-MCI dataset with a large number of change masks and captions of changes. Extensive experiments demonstrate the effectiveness of the proposed MCI model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes. To facilitate future research, we will make our dataset and codebase of the MCI model and Change-Agent publicly available at https://github.com/Chen-Yang-Liu/Change-Agent
@@ I Introduction
Current RSICI technology primarily encompasses change detection and change captioning. Change detection accurately localizes the spatial location of surface changes at the pixel level [6, 7, 8, 9, 10]. In contrast, change captioning aims at articulating the attributes and meanings of changes using natural language, emphasizing the semantic-level understanding of changes [11, 12, 13, 14, 15]. Despite significant advancements in both areas, their respective limitations prevent us from obtaining comprehensive change interpretation information by utilizing a single technology. Specifically, while change detection can accurately localize the changed area, it lacks a deeper understanding of changes’ underlying meanings, such as the characteristics of ground objects or the spatial relationships between them. Conversely, change captioning can provide rich semantic-level interpretation information but may fall short in providing precise pixel-level change localization. Therefore, there is a pressing need to explore a multi-level change interpretation (MCI) approach providing both pixel-level and semantic-level change information, facilitating precise change localization while delving into the essence and implications of the changes.
@@
Figure 1: The comparison between previous single technology and our Change-Agent. Our Change-Agent can simultaneously achieve precise pixel-level change detection and semantic-level change captioning. Besides, it boasts interactive capabilities, enabling users to communicate their queries or requirements regarding surface changes.
@@
Recent advancements in LLMs such as ChatGPT [16] and Llama2 [17] have been remarkable. Trained on vast corpora, LLMs have acquired extensive knowledge and demonstrated powerful capabilities in instruction comprehension, planning, reasoning, question answering, and text generation [18, 19]. These notable progress have provided robust support for our LLM-based Change-Agent. In this work, the LLM serves as the brain of the Change-Agent, assuming a pivotal role in internal scheduling, planning, and user-intent understanding. Integration with LLMs imbues our agent with enhanced flexibility and intelligence, enabling a nuanced understanding of user intentions and provision of customized change interpretation and intelligent analysis services such as changed object counting, estimation of change causes, prediction of future changes, etc.
Our Change-Agent facilitates accurate and comprehensive change interpretation of RS images. It reduces the workload and time costs for researchers while enhancing the efficiency and convenience of interaction between users and remote sensing data. Extensive experiments demonstrate the effectiveness of the proposed MCI model and highlight the promising potential of our Change-Agent. Besides, the integration of the MCI model and LLM presents novel perspectives for analyzing surface changes.
@@
  • We built a multi-level change interpretation dataset. The dataset contains bi-temporal images as well as diverse change detection masks and descriptive sentences. It provides a crucial data foundation for exploring multi-task learning for change detection and change captioning.
  • We propose a dual-branch MCI model, which can provide pixel-level and semantic-level change interpretation. Besides, we propose BI3 layers with LPE and GDFA to enhance the change interpretation capability of the model. Experiments validate the effectiveness of our method.
  • Leveraging the MCI model and an LLM, we construct a Change-Agent that achieves an interactive and comprehensive interpretation and analysis of surface changes. It has intelligent dialogue and customized service capabilities, opening up new opportunities for intelligent remote sensing applications.
@@ II Related work
A comprehensive change interpretation model should provide both pixel-level and semantic-level change information. However, previous work has primarily focused on either change detection or change captioning alone. Change detection aims to generate pixel-level change maps revealing the locations of changes in bi-temporal images, while change captioning aims to describe the changes via language. In this section, we will briefly review these approaches.
@@ II-A Remote Sensing Change Detection
Early change detection approaches were traditionally categorized into algebra-based, transformation-based, and classification-based methods. These methods relied on techniques such as change vector analysis (CVA) [20, 21], principal component analysis (PCA) [22, 23], multivariate alteration detection (MAD) [24, 25], and post-classification comparisons [26, 27]. Traditional change detection methods have laid the foundation for modern techniques in remote sensing change detection [28].
With the emergence of deep learning, the field of remote sensing change detection has witnessed significant advancements [29, 30, 31, 32, 33]. Compared to traditional methods, deep learning significantly improves the performance of change detection with its superior multi-temporal feature learning capability. These methods utilize various architectures including convolutional neural networks (CNNs) [34, 35, 36, 37], recurrent neural networks (RNNs) [38, 39, 40, 41, 42], auto-encoders [43, 44, 45], and Transformers [46, 47, 6, 48, 49, 50]. For instance, Chen et al. [46] introduced a bi-temporal image Transformer (BIT) [46], where images are represented as semantic tokens, and the Transformer encoder and decoder refine visual tokens to identify changed areas efficiently and effectively. Li et al. [6] proposed an end-to-end TransUNetCD model based on Transformer and UNet structure. The model utilizes a Transformer encoder to extract global contextual features. A CNN-based decoder upsamples the multi-scale features to generate the change map. Bandara et al. proposed a pure transformer model named ChangeFormer [51], whose backbone is a siamese SegFormer [52]. Zhang et al. [9] propose a deep siamese network with multi-scale multi-attention. They enhance global representation using a contextual attention module and fuse multi-scale features from the siamese feature extractor for detecting objects of varying sizes and irregularities. Some surveys, such as [2, 1, 5] provide a more comprehensive review of change detection in deep learning.
@@ II-B Remote Sensing Change Captioning
Remote sensing change captioning (RSCC) is a recently emerged multimodal task involving remote sensing image processing and natural language generation. Given the intricacy of this task, prevailing methodologies predominantly leverage deep learning techniques. Early methods in this domain largely drew inspiration from the encoder-decoder framework commonly utilized in image captioning tasks. Within this framework, the encoder is responsible for extracting features from bi-temporal images and performing the difference-aware fusion of these features. The decoder typically employs recurrent neural networks or Transformers to convert visual features into natural language. Compared to tasks like image captioning, which extract semantic representations from single images, change description faces the significant challenge of extracting robust semantic change representations.
Chouaf and Hoxha et al. [11, 12] pioneered remote sensing change captioning by introducing a model that employs a pre-trained CNNs [53] as the encoder and tried to utilize an RNN and Support Vector Machines (SVMs) as the decoder to generate textual descriptions. To improve the change-awareness of the model, they proposed an image-based fusion strategy and a feature-based fusion strategy. To facilitate research on change captioning, Liu et al. [13] introduced a large LEVIR-CC dataset and benchmarked some methods from other fields. Besides, they proposed a Transformer-based method named RSICCformer. The approach incorporates Siamese cross-encoding modules and multistage bitemporal fusion modules, emphasizing change regions through differencing features and capturing multiple changes of interest.
In a further advancement, some attention-based improvements have emerged [54, 14, 55]. For instance, PSNet [54] enhances the perception of changed objects of varying sizes using a progressive scale-aware network with difference perception layers and scale-aware reinforcement modules. Chang et al. [14] proposed an attention network containing a hierarchical self-attention module followed by a residual unit for identifying change-related attributes and constructing the semantic change embedding. Different from the encoder-decoder framework, Liu et al. [15] proposed a new paradigm to decouple the RSCC task by designing an image-level classifier and a feature-level encoder. Besides, a multi-prompt learning strategy is proposed to effectively exploit the potential of the LLM for captioning.
@@ III LEVIR-MCI Dataset
The proposed LEVIR-MCI dataset is an extension of our previously established change captioning dataset, LEVIR-CC [13]. In this dataset, we further provided each pair of bi-temporal images with additional change detection masks highlighting changed roads and changed buildings. Unlike its predecessor, the LEVIR-MCI dataset provides diverse annotations from different interpretation perspectives for each pair of images, further enhancing its utility for comprehensive change interpretation.
@@
Figure 3: Distribution of the scale and deformation of changed roads and buildings. The ’area’ on the horizontal axis represents the area of a single object, and the ’bbox_area’ on the vertical axis represents the area of the corresponding rectangular bounding box. The dispersion of points offers insights into the diversity of object scale and deformation.
@@ III-B Scale and Deformation of Changed Objects
An insightful examination of object scale and deformation is conducted through an analysis of road and building areas and the corresponding rectangular bounding boxes, as illustrated in Fig. 3. The ’area’ on the horizontal axis represents the area of a single object, and the ’bbox_area’ on the vertical axis represents the area of the corresponding bounding box. The dispersion of points offers insights into the diversity of object scale and deformation. Notably, points of roads exhibit a relatively broader spectrum with dispersed diversity, attributable to their narrow and curved nature. In contrast, points of buildings display a more concentrated distribution, reflecting their predominantly rectangular form. This analysis elucidates the intricate interplay between object scale and shape, shedding light on the nuanced characteristics of changed objects.
Figure 4: The overview of Change-Agent. The Change-Agent is equipped with an MCI model and an LLM, serving as its eyes and brain, respectively. The proposed LEVIR-MCI dataset provides a data foundation for training the MCI model.
@@ IV-B Bi-temporal Iterative Interaction Layer
Motivated by the analysis presented in Section III-B regarding the various scales and deformations of roads and buildings, we propose the LPE module with a residual structure. The LPE module employs convolution kernels of different sizes to extract multiple feature maps across different scales. This design enriches the diversity of feature information and improves the model’s local feature perception capabilities. Assuming the bi-temporal features are denoted as ${X_{T_{1}}}\in\mathbb{R}^{N\times C}$ and ${X_{T_{2}}}\in\mathbb{R}^{N\times C}$, the LPE module can be formulated as follows:
@@
The GDFA module leverages differencing features to generate spatial attention weights and perform the interaction and fusion between features. This facilitates the model to focus on changes of interest and ignore irrelevant disturbances. Taking the left GDFA module in Figure 5 as an example, assuming $X_{lp_{1}}\in\mathbb{R}^{N\times C}$ and $X_{lp_{2}}\in\mathbb{R}^{N\times C}$ are the bi-temporal features obtained from the LPE module, the GDFA module can be formulated as follows:
@@
Figure 7: The structure of the change captioning branch. The domain bridging module further bridges the conversion from the visual domain to the textual domain. The Transformer decoder generates captions.
@@ IV-C Change Detection and Captioning Branch
The structure of the change detection branch is depicted in Fig. 6. Leveraging multi-scale bi-temporal features extracted from the backbone network, this branch facilitates refined mask prediction. Specifically, multiple Bi-temporal Iterative Interaction (BI3) layers with residual connections iteratively enhance and refine bi-temporal high-level features, effectively capturing semantic changes. Since the low-level features contain more detailed information, they are crucial for refined change boundary detection. We further incorporate multiple Convolution-Based Bi-temporal Fusion (CBF) modules for bi-temporal feature fusion across four scales. Subsequently, through deconvolution, features are progressively integrated from bottom to top, bolstering the model’s discrimination capability for changes and enhancing change detection accuracy. For the input $X_{T_{1}}^{s},X_{T_{2}}^{s}(s=1,2,3,4)$, the CBF module can be represented as follows:
@@
The change captioning branch focuses on semantic-level change interpretation, translating visual changes into textual descriptions, as depicted in Fig. 7. This branch leverages multiple BI3 layers to interactively process high-level semantic features extracted from the backbone network, thereby obtaining bi-temporal visual features revealing changes of interest. Subsequently, a convolution-based domain bridging module further processes the bi-temporal features to facilitate the transition from the visual domain to the textual domain. Finally, these processed features are fed into a Transformer decoder to generate descriptive sentences elucidating the changes. The processing flow of the domain bridging module can be expressed as follows:
$\mathrm{\Phi_{DB}}(X_{T_{1}},X_{T_{2}}))=F_{1}+\mathrm{conv_{1\times 1}}(F_{2})$ (15)

### v3

@@ Abstract
Monitoring changes in the Earth’s surface is crucial for understanding natural processes and human impacts, necessitating precise and comprehensive interpretation methodologies. Remote sensing satellite imagery offers a unique perspective for monitoring these changes, leading to the emergence of remote sensing image change interpretation (RSICI) as a significant research focus. Current RSICI technology encompasses change detection and change captioning, each with its limitations in providing comprehensive interpretation. To address this, we propose an interactive Change-Agent, which can follow user instructions to achieve comprehensive change interpretation and insightful analysis, such as change detection and change captioning, change object counting, change cause analysis, etc. The Change-Agent integrates a multi-level change interpretation (MCI) model as the eyes and a large language model (LLM) as the brain. The MCI model contains two branches of pixel-level change detection and semantic-level change captioning, in which the BI-temporal Iterative Interaction (BI3) layer is proposed to enhance the model’s discriminative feature representation capabilities. To support the training of the MCI model, we build the LEVIR-MCI dataset with a large number of change masks and captions of changes. Experiments demonstrate the SOTA performance of the MCI model in achieving both change detection and change description simultaneously, and highlight the promising application value of our Change-Agent in facilitating comprehensive interpretation of surface changes, which opens up a new avenue for intelligent remote sensing applications. To facilitate future research, we will make our dataset and codebase publicly available at https://github.com/Chen-Yang-Liu/Change-Agent
@@ I Introduction
Current RSICI technology primarily encompasses change detection and change captioning. Change detection accurately localizes the spatial location of surface changes at the pixel level [6, 7]. In contrast, change captioning aims at articulating the attributes and meanings of changes using natural language, emphasizing the semantic-level understanding of changes [8, 9, 10, 11, 12]. Despite significant advancements in both areas, their respective limitations prevent us from obtaining comprehensive change interpretation information by utilizing a single technology. Specifically, although change detection can accurately localize the changed area, it lacks a deeper semantic-level understanding of changes’ underlying meanings, such as the characteristics of ground objects (e.g., category, color, shape), the spatial relationships between them (e.g., “beside”, “around”) and dynamics of changes (e.g., “appear”, “removed”). Conversely, change captioning can provide rich semantic-level interpretation information but may fall short in providing precise pixel-level change localization. Therefore, there is a pressing need to explore a multi-level change interpretation (MCI) approach providing both pixel-level and semantic-level change information, facilitating precise change localization while delving into the essence and implications of the changes.
@@
Figure 1: The comparison between previous single technology and our Change-Agent. Our Change-Agent can simultaneously achieve precise pixel-level change detection and semantic-level change captioning. Besides, it boasts interactive capabilities, enabling users to communicate their queries on surface changes.
@@
Recent advancements in LLMs such as ChatGPT [13] and Llama2 [14] have been remarkable. Trained on vast corpora, LLMs have acquired extensive knowledge and demonstrated powerful capabilities in instruction comprehension, planning, reasoning, question answering, and text generation [15, 16]. These notable progress have provided robust support for our LLM-based Change-Agent. In this work, the LLM serves as the brain of the Change-Agent, assuming a pivotal role in internal scheduling, planning, and user-intent understanding. Integration with LLMs imbues our agent with enhanced flexibility and intelligence, enabling a nuanced understanding of user intentions and provision of customized change interpretation and intelligent analysis services such as changed object counting, estimation of causes, prediction of changes, etc.
Our Change-Agent facilitates accurate and comprehensive change interpretation of RS images, offering novel perspectives for analyzing changes on the Earth’s surface. It reduces the workload and time costs for researchers while enhancing the efficiency and convenience of interaction between users and remote sensing data. Extensive experiments demonstrate the effectiveness of the proposed MCI model and highlight the promising application value of our Change-Agent.
@@
  • We develop an interactive Change-Agent, which uses external tools to achieve comprehensive interpretation and analysis of surface changes according to the user’s instructions. It has intelligent dialogue and customized service capabilities, opening a new avenue for intelligent remote-sensing applications.
  • We propose a multi-task learning-based MCI model with dual branches for change detection and change captioning. This model can simultaneously provide pixel-level and semantic-level change interpretation information, serving as the perception tool of the Change-Agent.
  • We built an MCI dataset named LEVIR-MCI, which contains bi-temporal images, corresponding change detection masks and descriptive sentences. The dataset provides a crucial data foundation for exploring multi-task learning for change detection and change captioning.
@@ II Related work
A comprehensive change interpretation model should provide both pixel-level and semantic-level change information. However, previous work has primarily focused on either change detection or change captioning alone. Change detection aims to generate pixel-level change maps revealing the locations of changes in bi-temporal images, while change captioning aims to describe the changes via language.
@@ II-A Remote Sensing Change Detection
Early change detection approaches were traditionally categorized into algebra-based, transformation-based, and classification-based methods. These methods relied on techniques such as change vector analysis (CVA) [17, 18], principal component analysis (PCA) [19, 20], multivariate alteration detection (MAD) [21, 22], and post-classification comparisons [23, 24]. Traditional methods have laid the foundation for modern change detection techniques [25].
With the emergence of deep learning[26, 27, 28, 29, 30, 31], the field of remote sensing change detection has witnessed significant advancements [32, 33, 34, 35, 36, 37, 38, 39]. Compared to traditional methods, deep learning significantly improves the performance of change detection with its superior multi-temporal feature learning capability. These methods utilize various architectures including convolutional neural networks (CNNs) [40, 2, 41], recurrent neural networks (RNNs) [42, 43, 44, 45, 46], auto-encoders [47, 48, 49], and Transformers [50, 51, 52, 4, 53, 54]. For instance, Chen et al. [50] introduced a bi-temporal image Transformer (BIT) [50], where images are represented as semantic tokens, and the Transformer encoder and decoder refine visual tokens to identify changed areas efficiently and effectively. Li et al. [52] proposed an end-to-end TransUNetCD model based on Transformer and UNet structure. The model utilizes a Transformer encoder to extract global contextual features. A CNN-based decoder upsamples the multi-scale features to generate the change map. Bandara et al. proposed a pure transformer model named ChangeFormer [55], whose backbone is a siamese SegFormer [56]. Zhang et al. [57] propose a deep siamese network with multi-scale multi-attention. They enhance global representation using a contextual attention module and fuse multi-scale features from the siamese feature extractor for detecting objects of varying sizes and irregularities. Compared to binary change detection, semantic change detection considers the semantic information of change categories, which is presented in the form of semantic masks. Some works, such as [58, 59], have significantly advanced the development of semantic change detection. However, higher-level semantic information, such as the relationships between changed objects, their colors, and spatial distribution, remains challenging to directly obtain. Some surveys, such as [60, 5] provide a more comprehensive review of change detection in deep learning.
@@ II-B Remote Sensing Change Captioning
Remote sensing change captioning (RSCC) is a recently emerged multimodal task involving remote sensing image processing and natural language generation. Given the intricacy of this task, prevailing methodologies predominantly leverage deep learning techniques. Early methods in this domain largely drew inspiration from the encoder-decoder framework commonly utilized in the image captioning task [61, 62, 63]. Within this framework, the encoder is responsible for extracting features from bi-temporal images and performing the difference-aware fusion of these features. The decoder typically employs recurrent neural networks or Transformers to convert visual features into natural language. Compared to tasks like image captioning, which extract semantic representations from single images, change description faces the significant challenge of extracting robust semantic change representations.
Chouaf and Hoxha et al. [8, 9] pioneered remote sensing change captioning by introducing a model that employs a pre-trained CNNs as the encoder and tried to utilize an RNN and Support Vector Machines (SVMs) as the decoder to generate textual descriptions. To improve the change-awareness of the model, they proposed an image-based fusion strategy and a feature-based fusion strategy. To facilitate research on change captioning, Liu et al. [10] introduced a large LEVIR-CC dataset and benchmarked some methods from other fields [64, 65]. Besides, they proposed a Transformer-based method named RSICCformer. The approach incorporates Siamese cross-encoding modules and multistage bitemporal fusion modules, emphasizing change regions through differencing features and capturing multiple changes of interest.
In a further advancement, some attention-based improvements have emerged [11, 66, 67, 68, 69, 70]. For instance, PSNet [68] enhances the perception of changed objects of varying sizes using a progressive scale-aware network with difference perception layers and scale-aware reinforcement modules. Chang et al. [11] proposed an attention network containing a hierarchical self-attention module followed by a residual unit for identifying change-related attributes and constructing the semantic change embedding. Unlike the encoder-decoder framework, Liu et al. [12] proposed a decoupling paradigm by designing an image-level classifier and a feature-level encoder. Besides, a multi-prompt learning strategy is proposed to effectively exploit an LLM for captioning.
Figure 2: The comparison of the proposed LEVIR-MCI dataset and the previous LEVIR-CC dataset. The LEVIR-MCI dataset is an extension of LEVIR-CC.
@@ III LEVIR-MCI Dataset
The proposed LEVIR-MCI dataset is an extension of our previously established change captioning dataset, LEVIR-CC [10]. As shown in Fig. 2, we further provided each pair of bi-temporal images with additional change detection masks highlighting changed roads and changed buildings. Unlike its predecessor, the LEVIR-MCI dataset provides diverse annotations from different interpretation perspectives for each pair of images, further enhancing its utility for comprehensive change interpretation.
@@
Figure 4: Distribution of the scale and deformation of changed roads and buildings. The “area” on the horizontal axis represents the area of a single object, and the “bbox_area” on the vertical axis represents the area of the corresponding rectangular bounding box. The value of the color bins represents the number of object instances. The dispersion of points offers insights into the diversity of object scale and deformation.
@@ III-B Scale and Deformation of Changed Objects
An insightful examination of object scale and deformation is conducted through an analysis of road and building areas and the corresponding rectangular bounding boxes, as illustrated in Fig. 4. The “area” on the horizontal axis represents the area of a single object, and the “bbox_area” on the vertical axis represents the area of the corresponding bounding box. The dispersion of points offers insights into the diversity of object scale and deformation. Notably, points of roads exhibit a relatively broader spectrum with dispersed diversity, attributable to their narrow and curved nature. In contrast, points of buildings display a more concentrated distribution, reflecting their predominantly rectangular form. This analysis elucidates the intricate interplay between object scale and shape, shedding light on the nuanced characteristics of changed objects.
Figure 5: The overview of Change-Agent is shown in (a). The Change-Agent is equipped with an MCI model and an LLM, serving as its eyes and brain, respectively. The proposed LEVIR-MCI dataset provides a data foundation for training the MCI model. (b) shows the overall structure of the MCI model.
@@ IV-B Bi-temporal Iterative Interaction Layer
Motivated by the analysis presented in Section III-B regarding the various scales and deformations of roads and buildings, we propose the LPE module with a residual structure. The LPE module employs convolution kernels of different sizes to extract multiple feature maps across different scales. This design enriches the diversity of feature information and improves the model’s local feature perception capabilities. Assuming the bi-temporal features are denoted as ${X_{T_{1}}}\in\mathbb{R}^{N\times C}$ and ${X_{T_{2}}}\in\mathbb{R}^{N\times C}$, the LPE module $\mathrm{\Phi_{LPE}}(\cdot)$ can be formulated as follows:
@@
The GDFA module leverages differencing features to generate spatial attention weights and perform the interaction and fusion between features. This facilitates the model to focus on changes of interest and ignore irrelevant disturbances. Taking the left GDFA module in Fig. 6 as an example, assuming $X_{lp_{1}}\in\mathbb{R}^{N\times C}$ and $X_{lp_{2}}\in\mathbb{R}^{N\times C}$ are the bi-temporal features obtained from the LPE module, the GDFA module $\mathrm{\Phi_{GDFA}}(\cdot)$ can be formulated as follows:
@@
Figure 8: The structure of the change captioning branch. The projection layer further bridges the conversion from the visual domain to the textual domain. The Transformer decoder generates captions.
@@ IV-C Change Detection and Captioning Branch
The structure of the change detection branch is depicted in Fig. 7. Leveraging multi-scale bi-temporal features extracted from the backbone network, this branch facilitates refined mask prediction. Specifically, multiple Bi-temporal Iterative Interaction (BI3) layers with residual connections iteratively enhance and refine bi-temporal high-level features, effectively capturing semantic changes. Since the low-level features contain more detailed information, they are crucial for refined change boundary detection. We further incorporate multiple Convolution-Based Bi-temporal Fusion (CBF) modules for bi-temporal feature fusion across four scales. Subsequently, through deconvolution, features are progressively integrated from bottom to top, bolstering the model’s discrimination capability for changes and enhancing change detection accuracy. For the input $X_{T_{1}}^{s},X_{T_{2}}^{s}(s=1,2,3,4)$, the CBF module $\mathrm{\Phi_{CBF}}(\cdot)$ can be represented as follows:
@@
The change captioning branch focuses on semantic-level change interpretation, translating visual changes into textual descriptions, as depicted in Fig. 8. This branch leverages multiple BI3 layers to interactively process high-level semantic features extracted from the backbone network, thereby obtaining bi-temporal visual features revealing changes of interest. Subsequently, a convolution-based projection layer further processes the bi-temporal features to facilitate the transition from the visual domain to the textual domain. Finally, these processed features are fed into a Transformer decoder to generate descriptive sentences elucidating the changes. The processing flow of the projection layer $\mathrm{\Phi_{Pr}}(\cdot)$ can be expressed as follows:
$\mathrm{\Phi_{Pr}}(X_{T_{1}},X_{T_{2}}))=F_{1}+\mathrm{conv_{1\times 1}}(F_{2})$ (15)